Solve the equations.
9p2-15p-6=0

Answer by stanbon(75887)

Solve the equations.
9p2-15p-6=0
----
Divide thru by 3 to get:
3p^2 - 5p - 2 = 0
---
3p^2-6p+p-2 = 0
---
3p(p-2)+(p-2) = 0
---
(p-2)(3p+1) = 0
p = 2 or p = -1/3
==========================
Cheers,
Stan H.
==========================